OpenStudy (anonymous):

A tank is full of water. Find the work required to pump the water out of the spout. (Use 9.8 m/s2 for g. Use 1000 kg/m3 as the density of water. Assume r = 9 m and h = 3 m.)

OpenStudy (amistre64):

missing some vital information such as shape of the tank and ... well i spose it said full of water and not half or third full

OpenStudy (anonymous):

sorry spherical

OpenStudy (amistre64):

murmur ... couldnt be cylindar could it lol; so what does the h refer to then?

OpenStudy (amistre64):

that clears things up

OpenStudy (amistre64):

we should determine what a typical slice of volume looks like then

OpenStudy (amistre64):



OpenStudy (amistre64):

i wonder if we want to use trigs in this ... cant tell if itd make it simpler or not

OpenStudy (anonymous):

i dont think it will make it simpler

OpenStudy (amistre64):

well, lets define the radius of a disc in terms of y then R = sqrt(9^2 - (9-y)^2) = sqrt(81 - (81 -18y +y^2)) = sqrt(81 - 81 +18y -y^2) = sqrt(18y -y^2) seems to work for the top half

OpenStudy (amistre64):

whereas: R = 9 sin(angle) seems pretty simple

OpenStudy (anonymous):

whats the angle

OpenStudy (amistre64):